# boxparampadding=1.5pt, padding-left=2pt, padding-right=2pt, height=7.5pt, background-color=lightgray, border-radius=2pt, border-right-width=1pt, border-bottom-width=1pt, border-color=gray ExploreSelf: Fostering User-driven Exploration and Reflection on Personal Challenges with Adaptive Guidance by Large Language Models

### v2

@@ A.2. Question Generation
¡initial_information/¿: Client’s initial brief introductory of difficulty, and the client’s background.
¡previous_session_log¿: Logs of sessions before the current session.
¡theme_of_session/¿: Theme of the current session.
A.3. Keyword Generation
@@
¡initial_information/¿: Client’s initial brief introductory of difficulty, and the client’s background.
¡previous_session_log¿: Logs of sessions before the current session.
¡question/¿: Question that the user is trying to response now.
A.4. Comments Generation
@@
Task: While the user is responding to the Socratic question, your task is to provide useful comments based on the context of user’s response. For example, if the user hasn’t started answering the question or is in very early phase, you might a) provide tips on how to approach the question. You might also provide b) encouraging, affirmation, supportive feedback if you find that the , in a user-tailored manner. It also could be the c) sub-question in answering the given question, or some d) insightful comment, or e) any other type of comment that might support user in responding to the question.
@@
¡initial_information/¿: Client’s initial brief introductory of difficulty, and the client’s background.
¡previous_session_log¿: Previous Logs of the session before the current session.
¡question/¿: Question that the user is trying to think about now.
¡current_response¿: Current response status of the user
A.5. Summary Generation
@@
Appendix B Statements of Pathway Subscale of Adult State Hope Scale

### v3

@@ B.2. Question Generation
<initial_information/>: Client’s initial brief introductory of difficulty, and the client’s background.
<previous_session_log>: Logs of sessions before the current session.
<theme_of_session/>: Theme of the current session.
B.3. Keyword Generation
@@
<initial_information/>: Client’s initial brief introductory of difficulty, and the client’s background.
<previous_session_log>: Logs of sessions before the current session.
<question/>: Question that the user is trying to response now.
B.4. Comments Generation
@@
Task: While the user is responding to the Socratic question, your task is to provide useful comments based on the context of user’s response. For example, if the user hasn’t started answering the question or is in very early phase, you might a) provide tips on how to approach the question. You might also provide b) encouraging, affirmation, or supportive feedback in a user-tailored manner. It also could be the c) sub-question in answering the given question, or some d) insightful comment, or e) any other type of comment that might support user in responding to the question.
@@
<initial_information/>: Client’s initial brief introductory of difficulty, and the client’s background.
<previous_session_log>: Previous Logs of the session before the current session.
<question/>: Question that the user is trying to think about now.
<current_response>: Current response status of the user
B.5. Summary Generation
@@
Appendix C Statements of Pathway Subscale of Adult State Hope Scale